OpenStudy (anonymous):

Q1-I have twice as many 5c coins as 20c coins. The total value of all my 5c and 20 c coins is $4.50. How many 5c coins do I have? Q2-. I have 20 coins consisting of 5c and 20c coins. The total value of the coins is $2.05. How many of each kind do I have? Please answer, Thanks!

OpenStudy (jack1):

lets retry that, sorry, this time keep it in cents, not dollars (lats equations i posted in a mix of both... my bad Q1. 5x + 20y = 450 2x = y now try, should get more reasonable results this time ;P

OpenStudy (anonymous):

still a bit confused, maths isn't my favourite subject.

OpenStudy (jack1):

Q2. x = number of 5c coins y = number of 20c coins you have 20 coins total so :x + y = 20 5x + 20y = 205 again... simultaneous equations

OpenStudy (anonymous):

I see, thanks! So going to bring home that A, also thanks for showing me how to work it out.

OpenStudy (jack1):

sweet good luck dude